# CUQDS: Conformal Uncertainty Quantification under Distribution Shift for Trajectory Prediction

### v2

@@ 8             $f_{\theta}(x^{t}),\hat{\sigma}(x^{t})\leftarrow{\tt CUQDS}{}_{\theta,\omega}(x^{t})$
3.2.1 Gaussian Process
Gaussian process regression often acts as the surrogate model to estimate the output distribution of existing models Erlygin et al. (2023). It introduces additional loss term to guild the learning process of existing models and reduce the output uncertainty. Gaussian process regression assumes any combinations of data samples follows different join distribution and thus good at capturing the nonlinear relationship among time series data samples. In this study, we introduce to utilize the Gaussian process regression method to estimate the preliminary output distribution of existing trajectory prediction models.
We view the base model $f_{\theta}$ as a black box, whose design can be an existing neural network structure in the literature (examples include Zhou et al. (2022); Liu et al. (2021); Zhou et al. (2023); Liang et al. (2020); Zhong et al. (2022)). We aim to estimate the output distribution $(Y|x,\theta\sim\mathcal{N}\left(\hat{\mu}\left(x\right),\hat{{\sigma}}^{2}\left(x\right)\right))$ corresponding to the time series trajectory inputs $x$ by a learning-based Gaussian process regression module Rasmussen et al. (2006). We treat the output of base model $f_{\theta}\left(\cdot\right)$ on $x$ as the mean value $\hat{\mu}\left(x\right)$ of the estimated output distribution and the correspond variance $\hat{{\sigma}}^{2}\left(x\right)$ as the uncertainty of $f_{\theta}\left(x\right)$. In real-world settings of autonomous vehicles, the collected historical trajectories of the target vehicles inevitably contain noisy data due to the processing limitations during perception and object tracking steps. Such noise usually hardly captured by the trajectory prediction models such that greatly impact the output trajectory and the correspond uncertainty. To mitigate this problem, we model such noisy impact as $\epsilon\in\mathbb{R}^{D}$ and assume it follows Gaussian distribution $\mathcal{N}\left(0,{\sigma}_{\epsilon}^{2}\right)$ with zero mean for simplicity and variance of ${\sigma}_{\epsilon}^{2}\in\mathbb{R}^{D}$. ${\sigma}_{\epsilon}$ is set as trainable vector. Now the trajectory prediction process of base model is denoted as $Y|_{x}=\hat{\mu}(x)+\epsilon$, where $(x,y)$ is any data samples from $\mathcal{D}$. Then we have:
where $k(\cdot,\cdot)$ is the kernel function or the covariance function. For example, $k\left(x,x^{\prime}\right)$ quantifies the similarity between trajectory input data $x$ and $x^{\prime}$.
3.2.2 Learnable Kernel Function
@@
3.2.3 Eigenvector Inducing Variables
As shown in Eq. 3, the covariance $\hat{{\sigma}}^{2}(x)$ in standard Gaussian process regression method is typically optimized based on the whole training data. However, this can be computationally expensive when the number of training samples ($N_{1}$) is large, given that the entire Gaussian process regression model requires $O(N_{1}^{3})$ computational complexity and $O(N_{1}^{2})$ memory complexity. To alleviate such complexity while maintaining the effectiveness of the Gaussian process regression module, we propose a modification to approximate the covariance $\hat{{\sigma}}^{2}(x)$ in Eq. 3 by extracting $M\in\mathbb{R}$, $M\ll N_{1}$, inducing variables $\{v_{i}\}_{i=1}^{M}$, $v_{i}\in\mathbb{R}^{L\times D}$ by Principal Component Analysis (PCA). These $M$ inducing variables summarize the key information of training dataset. More specifically, we first standardize each input trajectory data sample $x_{i}\in\mathbb{R}^{L\times D}$ in training dataset $\mathcal{D}_{1}$ by its own mean and standardisation along time dimension. Then we build the $N_{1}\times N_{1}$ covariance matrix by $cov(x,x^{\prime})=\textbf{\em E}[(x-\textbf{\em E}(x))(x^{\prime}-\textbf{\em E}(x^{\prime}))]$. The further steps follow the standard PCA processes.
3.2.4 Inference under Distribution Shift
@@
3.3 Uncertainty Quantification under Distribution Shift through Calibration
The learning-based Gaussian process regression module is prone to overfit the training data and provide overconfident uncertainty estimation under distribution shift. Moreover, such uncertainty is more concerned about the aleatoric uncertainty, referring to the probabilistic nature of noise in data Kendall and Gal (2017), but fails to consider the performance limitation of base model on different trajectory data, namely the epistemic uncertainty. To solve these problems, we propose a conformal P control module to calibrate the output uncertainty from the Gaussian process regression module by considering the performance of base model under potential distribution shift. The idea of this method is based on both the conformal prediction Shafer and Vovk (2008) and the P control in Angelopoulos et al. (2023). However, different from standard conformal prediction, we assume the training data and testing data follow different distributions. We initialize the conformal quantile using the validation data and keep updating it after every prediction step in the inference stage to provide trustworthy uncertainty quantification.
3.3.1 Conformal Prediction for Long-run Coverage
We aim to design a statistical-based method to calibrate the preliminary output uncertainty from the Gaussian process regression module in the inference stage. Our goal is to achieve a long-run average coverage rate in time, ensuring that the calibrated uncertainty interval covers the true target trajectories with a probability of $1-\alpha$ Angelopoulos et al. (2023), namely $\sum_{t=1}^{T}\texttt{err}^{t}/T\rightarrow\alpha$ as $T\rightarrow\infty$. $\texttt{err}^{t}=\mathbbm{1}(y^{t}\notin\mathcal{C}^{t})$ and $\alpha\in(0,1)$ is a predefined target error rate threshold. In particular, we seek to design a conformal score function $s(x,y)$ (smaller scores encode better agreement between $x$ and $y$), construct a conformal prediction set $\mathcal{C}^{t}(y|s(x^{t},y)\leq q^{t})$, estimate the conformal quantile $\hat{q}^{t}$ from the updated prediction set, and then calibrate the estimated uncertainty interval from the Gaussian process regression module by the updated conformal quantile. We define the score function as
@@ 3.3.1 Conformal Prediction for Long-run Coverage
where $y$ is the ground truth, $f_{\theta}(x)$ and $\sigma(x)$ are respectively the mean and std of the output distribution of the base model. Then, the conformal prediction set at time $t$ equals to an uncertainty interval $\{y|y\in\left[f_{\theta}(x^{t})-q^{t}{\sigma}\left(x^{t}\right),f_{\theta}(x^{t})+q^{t}{\sigma}\left(x^{t}\right)\right]\}$. Now the key of uncertainty calibration under potential distribution shift is to update the conformal quantile $q^{t}$ by considering model performance on recent inputs.
3.3.2 Uncertainty Calibration in CUQDS Framework
To update the conformal quantile $q^{t}$ under distribution shift, we first initialize the conformal quantile using the validation data as shown in Alg.1. In particular, during each validation iteration, we first initialize the score set $S$ which records the conformal scores and the error set $E$ which records the coverage errors of uncertainty intervals as empty. For each data sample $(x^{t},y^{t})$ in validation data $D_{2}$, we first compute the predicted trajectory $f_{\theta}(x^{t})$ from the base model and the standard deviation $\hat{\sigma}(x^{t})$ from the Gaussian process regression module for the trajectory input $x^{t}$. The conformal score for this prediction is calculated using the score function $s(x^{t},y^{t})$, which evaluates the model’s performance for current data sample. Then, a conformal prediction set $\mathcal{C}^{t}$ is computed to determine whether the true label $y^{t}$ falls within the estimated uncertainty interval. The error $e^{t}$ is set to 1 if $y^{t}$ is not within $\mathcal{C}^{t}$, indicating a prediction error, and 0 otherwise. The estimated conformal score $s(x^{t},y^{t})$ is added to the score set $S$, and the error $e^{t}$ is added to the error set $E$. The conformal quantile $\hat{q}^{t}$ is then updated based on both the existing conformal scores and coverage errors. The updating rule involves a learning rate $\eta=\beta\max(S)$ and adjusts $\hat{q}^{t}$ according to the formula
@@ 3.3.2 Uncertainty Calibration in CUQDS Framework
During the testing stage, the conformal P control module inherits the conformal quantile, the score set ${S}$, and the error set ${E}$ from the final validation iteration. The following conformal quantile updating steps are similar as in the validation stage as shown in Alg.2. For each time step $t$ in the testing stage, we get the predicted trajectory $f_{\theta}(\tilde{x}^{t})$ from the well trained base model and variance $\hat{\sigma}^{2}(\tilde{x}^{t})$ from the well trained Gaussian process regression module. We then calibrate the output uncertainty and construct the uncertainty interval by using the current conformal quantile $\hat{q}^{t}$,
@@
1
Input: the well-trained framework ${\tt CUQDS}{}_{\theta,\omega}$, the conformal quantile $\hat{q}^{1}$, the error set $E$, the conformal score set ${S}$, the error rate $\alpha$ from the training stage, score function $s(\cdot)$.
Output: predicted trajectory and the correspond calibrated uncertainty interval.
Data: testing data $\tilde{D}=\{\left(\tilde{x}^{t},\tilde{y}^{t}\right)\}_{t=L-1}^{T_{3}}$
@@
4 Experiment
4.1 Experimental Setups
Dataset & Key Setups: We use the Argoverse 1 motion forecasting dataset Chang et al. (2019) to verify the efficacy of our approach. This dataset collects trajectory data from Miami and Pittsburgh, with a sample rate of 10 Hz. Given that the ground truth future trajectories are not provided in this official test sequences but are essential to our CUQDS in the testing stage to update the conformal quantile $\hat{q}^{t}$, we repartition the sequences. In particular, we split the 205,942 official training sequences into 166,470 training data and 39,472 validation data, and use the official 39,472 validation sequences as testing data in this study.
@@ 4.1 Experimental Setups
Uncertainty Evaluation Metrics: To further verify the efficacy of our CUQDS in reducing the predicted uncertainty to provide narrow while accurate uncertainty interval, we adopt the Negative Log-Likelihood (NLL) Feng et al. (2021) to assess the level of uncertainty in the predicted distribution. Lower values connote a higher degree of precision in uncertainty estimation and narrower uncertainty interval.
4.2 Baselines
In this study, we employ the following models as baselines for comparison.
  1. 1.
    HiVT Zhou et al. (2022): HiVT is a transformer-based trajectory prediction model which models the output distribution as a mixture Laplace distribution. We set hidden units as 64 and radius as 50.
  2. 2.
    LaneGCN Liang et al. (2020): LaneGCN is a graph convolutional network based model that predicts the $k$ possible future trajectories and their confidence scores by a MLP-based prediction header.
  3. 3.
    LBA Zhong et al. (2022): LBA improves trajectory forecasting accuracy of existing models by fusing information from observed trajectories, HD maps, and local behavior data.
  4. 4.
    LBF Zhong et al. (2022): LBF improves trajectory forecasting accuracy of existing models by employing a local-behavior-free prediction framework which infers the impact of missing data when the local behavior historical data is insufficient or unavailable.
  5. 5.
    SPCI Xu and Xie (2023): SPCI proposes a learning-based estimator to predict the conditional quantile directly and bases on split conformal prediction to provide conformal prediction interval for existing time series forecasting models. We implement additional module whose structure is the same as the estimator to predict the variance of the predicted trajectory and add the KLD Meyer and Thakurdesai (2020) loss term to reduce the estimated uncertainty.
@@ 4.2 Baselines
4.3 Main Results
We adopt HiVT, LaneGCN, LBA, and LBF as base models and apply our CUQDS on them to verify the efficacy of CUQDS. To compare our CUQDS with state-of-the-art methods, we also apply SPCI into the above four base models. As presented in Table. 1 and Table. 2, our CUQDS improves the prediction accuracy 7.07% on average. Compared with incorporating SPCI in the base models, our CUQDS reduces the output uncertainty which quantified by NLL by 25.41% on average, and improves the average coverage rate of the estimated uncertainty interval by 21.02%. The results indicate that our CUQDS is capable of providing the trust worthy uncertainty quantification for the output trajectory of base model under potential distribution shift. Our estimated uncertainty intervals are narrow and achieve high coverage rate of covering the true target trajectories.
Compared with the transformer-based model HiVT, applying our CUQDS in HiVT provides the output uncertainty of HiVT by considering the distribution difference between the training and testing data. Compared with the base models of LaneGCN, LBA, and LBF who provide a confidence score for each output trajectory, our CUQDS provides the output distribution of base model instead of point estimates and calibrates the correspond uncertainty by taking into account the model performance on recent inputs. SPCI try to predict the conformal quantile by the learning-based estimator to provide the prediction interval. However, such estimator is prone to overfit the training data and invalid under distribution shift.
4.4 Ablation Study
$\bullet$ Conformal Prediction VS conformal P control module: We conduct ablation study on replacing the P control uncertainty calibration module with the standard split conformal prediction Shafer and Vovk (2008). More specificity, the standard split conformal prediction (CP) estimates the conformal quantile $\hat{q}$ as the $\frac{(1-\alpha)(N_{2}+1)}{N_{2}}$ smallest element in the conformal score set $\mathcal{S}={\{s\left(x_{i},y_{i}\right)\}}_{i=1}^{N_{2}}$ from validation dataset. This conformal quantile $\hat{q}$ is only calculated once and fixed in the testing stage to calibrate the uncertainty.
@@ 4.4 Ablation Study
$\bullet$ Gaussian process regression module VS direct modeling: To validate the effectiveness of our Gaussian process regression module, we replace the Gaussian process regression module with the self-attention based design Mao et al. (2023) to predict the variance of the output distribution and add the KLD loss term to reduce the estimated uncertainty during the training stage. We apply this setting on the base model HiVT. As shown in Table. 4, both of the methods achieve better prediction accuracy than without considering uncertainty in the base model. Our CUQDS slightly exceed the direct modeling method in all prediction accuracy evaluation metrics. The results prove that our CUQDS is sufficient in providing output uncertainty estimation for the trajectory prediction base model and improving the prediction accuracy.
5 Conclusion & Discussion

### v3

@@ 11             $f_{\theta}(x^{t}),\hat{\sigma}(x^{t})\leftarrow{\tt CUQDS}{}_{\theta,\omega}(x^{t})$
In this study, we introduce to utilize the Gaussian process regression method as introduced in preliminary to estimate the preliminary output distribution of existing trajectory prediction models. In our case, the base model $f_{\theta}$ can be any existing deep neural network structure in the literature, examples include (Zhou et al. 2022; Liu et al. 2021; Zhou et al. 2023; Liang et al. 2020; Zhong et al. 2022)). Moreover, in real-world settings of autonomous vehicles, the collected historical trajectories of the target vehicles inevitably contain noisy data during perception and object tracking steps. Such noise usually hardly captured by the trajectory prediction models such that greatly impacts the output trajectory and the corresponding uncertainty. To mitigate this problem, we model such noisy impact as $\epsilon\in\mathbb{R}^{D}$ as in the standard Gaussian process regression method.
Learnable Kernel Function
@@
Eigenvector Inducing Variables
As shown in Eq. 3, the covariance $\hat{{\sigma}}^{2}(x)$ of standard Gaussian process regression method is typically optimized based on the whole training data. However, this can be computationally expensive when the number of training samples ($N_{1}$) is large, given that the entire Gaussian process regression model requires $O(N_{1}^{3})$ computational complexity and $O(N_{1}^{2})$ memory complexity. To alleviate such complexity while maintaining the effectiveness of the Gaussian process regression module, we propose a modification to approximate the covariance $\hat{{\sigma}}^{2}(x)$ in Eq. 3 by extracting $M\in\mathbb{R}$, $M\ll N_{1}$, inducing variables $\{v_{i}\}_{i=1}^{M}$, $v_{i}\in\mathbb{R}^{L\times D}$ by Principal Component Analysis (PCA). These $M$ inducing variables summarize the key information of training dataset. More specifically, we first standardize each input trajectory data sample $x_{i}\in\mathbb{R}^{L\times D}$ in training dataset $\mathcal{D}_{1}$ by its own mean and standardization along time dimension. Then we build the $N_{1}\times N_{1}$ covariance matrix by $cov(x,x^{\prime})=\textbf{\em E}[(x-\textbf{\em E}(x))(x^{\prime}-\textbf{\em E}(x^{\prime}))]$. The further steps follow the standard PCA processes.
Inference under Distribution Shift
@@
Uncertainty Quantification under Distribution Shift through Calibration
The learning-based Gaussian process regression module is prone to overfit the training data and provide overconfident uncertainty estimation under distribution shift. To solve this, we propose a statistical-based conformal P control module to calibrate the output uncertainty from the Gaussian process regression module by considering the performance of base model under distribution shift. The idea of this method is based on both the conformal prediction (Shafer and Vovk 2008) and the P control in (Angelopoulos, Candes, and Tibshirani 2023). However, different from standard conformal prediction, we assume the training/validation data and testing data follow different distributions, and the distribution of testing data shifts over time. Instead of using the fixed conformal quantile $q$ as in standard conformal prediction, our proposed module updates the conformal quantile $q^{t}$ in time step $t$ based on model performance under distribution shift.
[…]
Uncertainty Calibration in CUQDS Framework
To update the conformal quantile $q^{t}$ under distribution shift, we first initialize the conformal quantile using the validation data as shown in Alg. 1. For each data sample $(x^{t},y^{t})$ in validation data $D_{2}$, we first compute the corresponding conformal score $s(x^{t},y^{t})$ and the conformal prediction set $\mathcal{C}^{t}$ as standard conformal prediction. We define the score function $s(x,y)$ to evaluate the model performance for each data sample
@@
where $y$ is the ground truth, $f_{\theta}(x)$ and $\sigma(x)$ are respectively the mean and std of the output distribution of the base model.
Moreover, we introduce an additional error set $E$ to record the coverage errors of uncertainty intervals. The error $e^{t}$ is set to 1 if $y^{t}$ is outside of $\mathcal{C}^{t}$, indicating a prediction error, and 0 otherwise. The estimated conformal score $s(x^{t},y^{t})$ is added to the score set $S$, and the error $e^{t}$ is added to the error set $E$. The conformal quantile $\hat{q}^{t}$ is then updated based on both the existing conformal scores and coverage errors. The updating rule involves a learning rate $\eta=\beta\max(S)$ and adjusts $\hat{q}^{t}$ according to the formula
@@
During the inference stage, the conformal P control module inherits the conformal quantile, the score set ${S}$, and the error set ${E}$ from the final validation iteration. The conformal quantile updating steps in the inference stage as shown in Alg. 2 are similar as in the validation stage. For each time step $t$ in the inference stage, we calibrate the output uncertainty and construct the uncertainty interval by using the updated conformal quantile $\hat{q}^{t}$,
@@
1 Input: the well-trained framework ${\tt CUQDS}{}_{\theta,\omega}$, the conformal quantile $\hat{q}^{1}$, the error set $E$, the conformal score set ${S}$, the error rate $\alpha$ from the training stage, score function $s(\cdot)$.
2 Data: testing data $\tilde{D}=\{\left(\tilde{x}^{t},\tilde{y}^{t}\right)\}_{t=L-1}^{T_{3}}$.
3 Output: predicted trajectory and the correspond calibrated uncertainty interval.
@@
Experiment
Experimental Setups
Dataset & Key Setups: We use the Argoverse 1 motion forecasting dataset (Chang et al. 2019) to verify the efficacy of our approach. This dataset collects trajectory data from Miami and Pittsburgh, with a sample rate of 10 Hz. Given that the ground truth future trajectories are not provided in this official test sequences but are essential to our CUQDS in the testing stage to update the conformal quantile $\hat{q}^{t}$, we repartition the sequences. In particular, we split the 205,942 official training sequences into 166,470 training data and 39,472 validation data, and use the official 39,472 validation sequences as testing data in this study.
@@
Uncertainty Evaluation Metrics: To further verify the efficacy of our CUQDS in reducing the predicted uncertainty to provide narrow while accurate uncertainty interval, we adopt the Negative Log-Likelihood (NLL) (Feng et al. 2021) to assess the level of uncertainty in the predicted distribution. Lower values connote a higher degree of precision in uncertainty estimation and narrower uncertainty interval.
@@
Main Results
We adopt HiVT (Liang et al. 2020), LaneGCN (Liang et al. 2020), LBA (Zhong et al. 2022), LBF (Zhong et al. 2022), and Trajectron++ (Salzmann et al. 2020) as base models and apply our CUQDS on them to verify the efficacy of CUQDS. To compare our CUQDS with state-of-the-art methods, we also apply SPCI (Xu and Xie 2023) into the above five base models to serve as the alternative of our CUQDS. In specifically, we implement additional module whose structure is the same as the estimator in SPCI to predict the variance of the predicted trajectory and add the KLD (Meyer and Thakurdesai 2020) loss term to reduce the estimated uncertainty. As presented in Table. 1 and Table. 2, our CUQDS improves the prediction accuracy 7.07% on average. Compared with incorporating SPCI in the base models, our CUQDS reduces the output uncertainty which quantified by NLL by 25.41% on average, and improves the average coverage rate of the estimated uncertainty interval by 21.02%. The results indicate that our CUQDS is capable of providing the trust worthy uncertainty quantification for the output trajectory of base model under potential distribution shift. Our estimated uncertainty intervals are narrow and achieve high coverage rate of covering the true target trajectories.
Compared with the transformer-based model HiVT, applying our CUQDS in HiVT provides the output uncertainty of HiVT by considering the distribution difference between the training and testing data. Compared with the base models of LaneGCN, LBA, LBF, and Trajectron++ who provide a confidence score or the estimated distribution for each output trajectory, our CUQDS provides the output distribution of base model instead of point estimates and calibrates the correspond uncertainty by taking into account the model performance on recent inputs. SPCI try to predict the conformal quantile by the learning-based estimator to provide the prediction interval. However, such estimator is prone to overfit the training data and invalid under distribution shift.
Ablation Study
$\bullet$ Conformal Prediction VS conformal P control module: We conduct ablation study on replacing the P control uncertainty calibration module with the standard split conformal prediction (Shafer and Vovk 2008). More specificity, the standard split conformal prediction (CP) estimates the conformal quantile $\hat{q}$ as the $\frac{(1-\alpha)(N_{2}+1)}{N_{2}}$ smallest element in the conformal score set $\mathcal{S}={\{s\left(x_{i},y_{i}\right)\}}_{i=1}^{N_{2}}$ from validation dataset. This conformal quantile $\hat{q}$ is only calculated once and fixed in the testing stage to calibrate the uncertainty.
@@
$\bullet$ Gaussian process regression module VS direct modeling: To validate the effectiveness of our Gaussian process regression module, we replace the Gaussian process regression module with the self-attention based design (Mao et al. 2023) to predict the variance of the output distribution and add the KLD loss term to reduce the estimated uncertainty during the training stage. We apply this setting on the base model HiVT. As shown in Table. 4, both of the methods achieve better prediction accuracy than without considering uncertainty in the base model. Our CUQDS slightly exceed the direct modeling method in all prediction accuracy evaluation metrics. The results prove that our CUQDS is sufficient in providing output uncertainty estimation for the trajectory prediction base model and improving the prediction accuracy.
Conclusion & Discussion
@@
Acknowledgments
The work is supported by the National Science Foundation under Grants CNS-2047354 (CAREER), and the New England University Transportation Center (NEUTC). Funding for the NEUTC Program is provided by the Office of Assistant Secretary for Research and Innovation (OST-R) of the United States Department of Transportation. The recommendations of this study are those of the authors and do not represent the views of NEUTC.